OpenStudy (anonymous):

Help! How do I solve this? (Equations algebra 1) 2x=2(x-3)

OpenStudy (anonymous):

easy divide both sides by 2

OpenStudy (anonymous):

Distribute the 2 to the x and 3 then you will have 2x and 6 and then add the 2x and the other 2x and then divide -6 by 4 and then its x= -1.5

jimthompson5910 (jim_thompson5910):

2x=2(x-3) 2x=2x-6 ... distribute 2x - 2x = 2x - 6 - 2x ... Subtract 2x from both sides. 0x = 0x - 6 0 = 0 - 6 0 = -6 this is a contradiction which is never true so there are no solutions